solve the system of equations by graphing:
2x+y=7
x+3y=6

When I tried solving this I put y=-2x-7...So that would make it -7/2x but I don't know how I would graph it...and then the second part I have no clue about.

Answer by mananth(16946)

2x+y=7........1
x+3y=6...........2
Find the x & y intercept points
equation1
when x= 0 y= 7 ( 0 , 7 ) 7
when y = 0 x= 3.5 ( 3.5 , 0 ) 3 1/2
Equation 2
when x= 0 y= 2 ( 0 , 2 ) 2
when y = 0 x= 6 ( 6 , 0 )



(3,1)